# DiffClean: Diffusion-based Makeup Removal for Accurate Age Estimation

### v1

@@ 4.1 Datasets
Training/Fine-tuning: We used the UTKFace [63] dataset to fine-tune SSRNet for age estimation with a training set of 15,364 images and a testing set of 6,701 images. The original set of 20K images was split into training and testing splits (70:30) after filtering out images with ages $\geq$70 yrs. We adopted a two-step approach for fine-tuning DiffClean for makeup removal. First, we used 300 images (200 training, 100 validation) from MT dataset [30] following the guidelines in [48]. We further refined the model by fine-tuning it additionally on 300 images (200 training, 100 validation) from the UTKFace dataset after digitally applying makeup using EleGANt [56] with the combined loss. Thus, our training set, $D_{tr}$ consisted of 600 images in total. We used five reference images [30] to simulate makeup styles while using EleGANt. Our focus is makeup removal which is a many-to-one task, i.e., translate from any makeup style to recover the clean non-makeup image. So, we used a few representative synthetic makeup styles and tested the robustness of our solution in real-world settings.
Evaluation: We tested our proposed DiffClean on both synthetic makeup and real-world makeup. For synthetic makeup, we sampled 2,556 images from the Flickr-Faces-HQ Dataset (FFHQ) [26] by sampling 284 images from each of the 9 age groups (excluding $\geq 70$ yrs.) using the age bins provided by FFHQ-Aging [37]. Next, we applied EleGANt to curate synthetic makeup dataset to assess the performance of our makeup removal framework. For real-world makeup, we used 3,000 images from the BeautyFace [55] dataset and 355 images from the LADN dataset [16] for testing generalizability of our method.
Figure 4: Makeup removal results generated by different methods. (First column:) Original images. (Second column:) Baseline outputs from CLIP2Protect [43] and DiffAM [48]. (Third column:) Our outputs from DiffClean with SSRNet and CLIP age loss. Our method significantly removes makeup (DiffAM retains makeup) without visual artifacts (CLIP2Protect changes hair color and facial features).
4.2 Metrics
@@
We evaluated the model for age restoration and identity retention. For age estimation, we report the Mean Absolute Error (MAE) between ground truth and predicted age values (lower is better); Age group accuracy to assess if the predicted age is correctly grouped into age bins (higher is better); Minor/Adult Accuracy to evaluate if the predicted age $\geq 18$ to classify as adult, otherwise, minor (higher is better). For identity verification, we report the True Match Rate (TMR) at a False Match Rate (FMR) of 0.01% and further show the ROC curves.
4.3 Implementation Details
@@
The SSRNet model was fine-tuned with batch size 50 with Adam optimizer with weight decay as 1e-4, cosine annealing scheduler and learning rate set to 1e-3 for 200 epochs with early stopping. The input image size is 64$\times$64. The makeup removal module takes the makeup image, $\textbf{I}_{m}$, of size 256$\times$256 as input, and uses 40 DDIM inversion steps with 6 DDIM sampling steps and (total time steps=80), and noise scale for latent image generation is controlled by cosine beta scheduler[35]. We used the Adam optimizer with a learning rate of 4e-3. The model is fine-tuned for 5 epochs. All experiments were carried out on a single NVIDIA A100 GPU. Hyperparameters in Eqn. 6 were obtained using grid search on the validation set.
4.4 Baselines
We compare DiffClean with 2 baselines: (1) CLIP2Protect [43], and (2) DiffAM [48]. To implement CLIP2Protect as a baseline, we used the available pretrained checkpoint with ‘no makeup’ text prompt and set the target identity to the source identity. For adversarial makeup transfer, we used single target identity and ‘Matte’ makeup style as it achieved highest success [43]. To implement DiffAM as a baseline (reproducing paper results), we needed to fine-tune the pretrained model provided by the authors [30] on MT dataset. For adversarial makeup transfer experiments, we used the checkpoint provided by the authors [48], and the default parameters corresponding to the target model, reference image, and target identity.
5 Results and Discussion
We present qualitative results for subjective evaluation and quantitative results for objective evaluation. We compare with the baselines in terms of age estimation and identity retention. See Fig. 4 for visualizing outputs of images before and after makeup removal and the corresponding outputs produced by the baselines.
5.1 Age estimation
@@
We report the results of the predicted age estimated by MiVOLO after makeup removal in Table 1. We compare the predicted age of ‘Makeup’ images simulated using EleGANt and compare with the age labels of the original FFHQ images (without makeup) and compute the minor vs adult accuracy (%), age group accuracy (%), and mean absolute error (MAE). In Table 1, our method DiffClean outperforms both baselines, CLIP2Protect and DiffAM, and can successfully restore the age cues disrupted due to makeup by lowering the MAE (to 5.71) and improving the overall age group prediction accuracy (to 37%). Our objective focused on minor age groups, and results indicate that we achieve superior minor vs adult age classification of 88.6%.
Figure 5: Examples of makeup removal on sample images from CelebA-HQ [25] and CACD [4] datasets while mitigating differential effects of makeup. Each column consists of pairs of images before makeup removal (left) and after makeup removal (right). In the first column, DiffClean results in lower predicted ages, while in the second column, DiffClean results in similar predicted ages, and finally, in the third column, DiffClean results in higher predicted ages than the original.
Makeup exhibits a differential effect on facial age. It may result in a decrease or increase in apparent age, and in some cases, may not affect perceived age. Our method does not arbitrarily reduce predicted age in all cases; instead, it strategically erases the makeup traces to restore the original age-specific cues. Fig. 5 clearly depicts the differential age restoration where the predicted age after makeup removal may be lower, equal, or higher than the estimated age on makeup images.
Table 1: Results of age estimation in terms of minor vs adult accuracy (%) $\uparrow$, age group accuracy (%) $\uparrow$ and mean abolsute error (MAE) $\downarrow$. Our method DiffClean outperforms baselines and improves overall age estimation with respect to makeup images.
Metrics Makeup images
CLIP2Protect
(‘no makeup’) prompt
DiffAM
DiffClean (Ours)
SSRNet age loss
DiffClean (Ours)
with CLIP age loss
Minor/Adult Acc (%) 82.8 83.8 85.5 87.8 88.6
Age group Acc (%) 34.4 33.3 34.5 36.8 37.0
MAE 6.56 6.54 6.32 5.76 5.71
@@
DiffClean (Ours-SSRNet) 75.8 ($\uparrow$) 84.2 ($\uparrow$)
DiffClean (Ours-CLIP) 75.5 ($\uparrow$) 82.4 ($\uparrow$)
@@
Figure 6: (Top): ROC curve with FaceNet. (Middle): ROC curve with MobileFace. (Bottom): Biometric matching in terms of TMR (%) @FMR = 0.01% (higher is better) with FaceNet and MobileFace matchers. Our DiffClean improves overall performance.
5.2 Identity verification
@@
We present the results of biometric identity verification in Fig. 6 using FaceNet and MobileFace face recognition (FR) models at FMR=0.01%. We compute the ‘Original’ performance by comparing the original FFHQ face images (without makeup) with images simulated using EleGANt (with makeup). Baseline performances are computed by comparing the original FFHQ face images with outputs of respective baselines (makeup removed images). Finally, we report the biometric matching of our proposed method by comparing the original FFHQ face images with outputs of DiffClean (makeup removed images). We observe that makeup removal improves face verification as initially hypothesized. The True Match Rate (TMR) improves by 1.3% with FaceNet and 5.1% with MobileFace compared to ‘Original’ performance, while TMR improves by 8.0% with FaceNet and 8.9% with MobileFace compared to the baselines. Note both baselines degrade identity cues.
5.3 Real-world makeup generalizability
Real-world examples only include image with makeup, but not the actual age label or the original image before makeup. So, it is necessary to evaluate the underestimation and overestimation in predicted ages after makeup removal. We used 3,000 makeup images from the BeautyFace [55] dataset and 355 images from the LADN [16] dataset. We tested DiffClean (with SSRNet and CLIP-based age loss) on these images, and used MiVOLO to predict the age before and after makeup removal; see Fig. 7. We observe that DiffClean lowers both overestimation and underestimation errors induced by makeup. Results indicate that our method reasonably generalizes across a broad spectrum of real-world makeup styles. Refer to Tables 2 and 3 for detailed statistics. We further report the confidence interval (CI) and margin of error on the BeautyFace dataset using the student’s t-distribution at a 95% confidence level. CI = [2.88, 3.09] with margin of error = 0.10 on the underestimation errors, and CI = [2.75, 2.89] with margin of error of 0.07 on the overestimation errors. This shows that our method can reduce the effects of makeup on estimated age by ±$\pm 3$ years on average.
@@ 5.3 Real-world makeup generalizability
Table 2: Statistics of predicted ages on the real-world images from the BeautyFace dataset before and after makeup removal.
@@
After makeup
removal
@@
Table 3: Statistics of predicted ages on the real-world images from the LADN dataset before and after makeup removal.
Before makeup
removal
After makeup
removal
$\mu$$\pm$ $\sigma$ $\mu$$\pm$ $\sigma$
Under
estimation
$\mu$$\pm$ $\sigma$
Over
estimation
$\mu$$\pm$ $\sigma$
24.64 $\pm$ 3.88
@@
Figure 7: Results of DiffClean on real-world makeup images from LADN [16] (left) and BeautyFace [55] (right) datasets. Our method is capable of reducing makeup induced overestimation errors in predicted age (indicated in white) in practical settings.
5.4 Ablation study
Training Dataset Diversity. We fine-tuned on 600 images (300 from MT dataset with smoky-eyes, flashy, retro, Korean and Japanese makeup styles + 300 from UTKFace with EleGANt-based makeup transfer with five reference styles and adaptive interpolated shading). To further assess training data diversity, we augmented our training set, $D_{tr}$, and report MAE on validation set, as follows: 4.42 ($\lvert D_{tr}\rvert=600$), 4.55 ($\lvert D_{tr}\rvert=1,200$) and 4.68 ($\lvert D_{tr}\rvert=1,600$). We observe that MAE is relatively stable across different training dataset scale.
Influence of each loss component. We report MAE ($\downarrow$) on the validation set by fine-tuning separate models with loss components added incrementally, as follows: 9.28 ($\mathcal{L}_{clip}$), 7.43 ( $\mathcal{L}_{clip}+\mathcal{L}_{id}$), 4.70 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}$), and 4.42 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}+\mathcal{L}_{SSRNet-age}$). Evidently, the combined losses improve the overall performance.
Demographic fairness analysis. We conducted demographic fairness analysis on FFHQ using DeepFace library [42]; see Table LABEL:Tab:fair. Note FFHQ has more White images than Black and Indian images, but a balanced gender distribution. We observe that DiffClean does not exhibit demographic disparity.
Inference time. We compared the inference time of makeup removal using our method with CLIP2Protect. We observed that our method is 15$\times$ faster; Ours: 2 secs. and CLIP2Protect: 30 secs.
Table 4: Demographic fairness analysis on FFHQ using our DiffClean SSRNet/CLIP-age loss models in terms of MAE. Values in parentheses indicate count after being processed by DeepFace.

### v2

@@ IV-A Datasets
Training/Fine-tuning: We used the UTKFace [UTKFace] dataset to fine-tune SSRNet for age estimation with a training set of 15,364 images and a testing set of 6,701 images. The original set of 20K images was split into training and testing splits (70:30) after filtering out images with ages $\geq$70 yrs. We adopted a two-step approach for fine-tuning DiffClean for makeup removal. First, we used 300 images (200 training, 100 validation) from MT dataset [li2018beautygan] following the guidelines in [sun2024diffam]. We further refined the model by additionally fine-tuning it on 300 images (200 training, 100 validation) from the UTKFace dataset after applying makeup using EleGANt [yang2022elegant] with the combined loss. Thus, our training set $D_{tr}$ consisted of a total of 600 images.
Evaluation: We tested our proposed DiffClean on both synthetic makeup (on FFHQ) and real-world makeup (LADN, BeautyFace, Makeup-Wild) images.
Synthetic makeup data generation: We sampled 2,556 images from the Flickr-Faces-HQ Dataset (FFHQ) [karras2019style] by sampling 284 images from each of the 9 age groups (excluding $\geq 70$ yrs.) using the age bins provided by FFHQ-Aging [FFHQ_Aging]. Next, we applied EleGANt [yang2022elegant] that needs reference images for makeup transfer. We used five reference images [li2018beautygan] to simulate varied makeup styles, representing lip, face and eye makeup; see Fig. 2. We note that there can be other makeup styles but the representative reference styles selected for data creation mimic real-world samples.
Real-world makeup data curation: We used 3,000 images from the BeautyFace [Beautyface] dataset, 355 images from the LADN dataset [ladn_2019], and 384 images from the Makeup-Wild dataset [jiang2020psgan] for testing generalizability of our method across variations in pose, illumination, accessories and occlusions.
IV-B Metrics
@@
We evaluated DiffClean for (1) age restoration using Mean Absolute Error (MAE) between ground truth (if available) and predicted age values (lower is better); Age group accuracy to assess if the predicted age is correctly grouped into age bins (higher is better); Minor/Adult Accuracy to evaluate if the predicted age $\geq 18$ to classify as adult, otherwise, minor (higher is better). We evaluated DiffClean for (2) identity verification by reporting the True Match Rate (TMR) at a False Match Rate (FMR) of 0.01% and illustrated the ROC curves. Finally, we computed (3) image quality of generated images using SSIM and PSNR.
IV-C Implementation Details
@@
The SSRNet model was fine-tuned with batch size 50 with Adam optimizer with weight decay as 1e-4, cosine annealing scheduler and learning rate set to 1e-3 for 200 epochs with early stopping. The input image size is $\times64$. The makeup removal module takes the makeup image, $\textbf{I}_{m}$, of size $\times256$ as input, and uses 40 DDIM inversion steps with 6 DDIM sampling steps and (total time steps=80), and noise scale for latent image generation is controlled by cosine beta scheduler [nichol2021improved]. We used the Adam optimizer with a learning rate of 4e-3. The model is fine-tuned for 5 epochs. All experiments were carried out on a single NVIDIA A100 GPU. Hyperparameters in Eqn. 6 were obtained using grid search on the validation set (see sensitivity analysis in Sec. V-E5). We perform age estimation using MiVOLO [kuprashevich2023MiVOLO] (while training uses SSRNet) and MobileFace [chen2018mobilefacenets] and FaceNet [schroff2015facenet] for identity verification (while training uses ArcFace IRSE50 [hu2018squeeze]).
IV-D Baselines
We compared DiffClean with 6 baselines (4 GAN-based and 2 diffusion-based models) : BeautyGAN [li2018beautygan], LADN [ladn_2019], PSGAN++ [PSGAN++_2021] CLIP2Protect [shamshad2023clip2protect], DiffAM [sun2024diffam], and MAD [ruan2025mad]. We randomly selected six non-makeup images as references for baselines following [DeBeauty_2025].
Figure 4: Examples of differential effects of age estimation before and after makeup removal. $1^{st}$ column: DiffClean causes lower predicted ages than with makeup, thus reducing overestimation errors; $2^{nd}$ column: DiffClean preserves original age when there is minimal or no makeup; $3^{rd}$ column: DiffClean causes higher predicted ages than with makeup, thus reducing underestimation errors. Results srongly support that DiffClean retains the age of original image if there is no makeup present.
V Results and Discussion
We present qualitative results for subjective evaluation (Figs. 346, and 7). We compare with the baselines in terms of age estimation (Tables I and III), identity verification (Fig. 5) and visual quality (Table II).
V-A Age estimation
@@
We report the results of the predicted age estimated by MiVOLO after makeup removal on FFHQ dataset in Table I. We compute the Original performance by comparing the clean FFHQ face images (without makeup) with images simulated using EleGANt (with makeup), and compute the minor vs. adult accuracy (%), age group accuracy (%), and mean absolute error (MAE). In Table I, our method DiffClean outperforms all baselines and can successfully restore age cues disrupted due to makeup by lowering the MAE (by 0.85) and improving the minor/adult prediction accuracy (by 5.1%). Our objective is focused on minor age groups and the results indicate that we achieve a superior minor vs. adult age classification of 88.6%.
We further examined the the differential effect of makeup on facial age. Makeup can either decrease or increase the apparent age and, in some cases, may not affect perceived age at all. So, we present some qualitative examples of how DiffClean handles such cases in Fig. 4. The first column shows pairs of images with higher predicted age before makeup removal and lower predicted age after makeup removal. This is the most challenging case which is the main focus of our work, i.e., how to mitigate overestimation errors due to makeup. The second column represents the benign cases when there is little to no makeup and DiffClean retains the age. The third column shows pairs of images with lower predicted age before makeup removal and higher predicted age after makeup removal, thus mitigating underestimation errors. Our method does not arbitrarily reduce or increase the predicted age in all cases. Instead, it strategically erases makeup traces to restore the original age-specific cues.
TABLE I: Results of age estimation on FFHQ in terms of minor vs adult accuracy (%) $\uparrow$, age group accuracy (%) $\uparrow$ and mean absolute error (MAE) $\downarrow$. Bolded values indicate best results while underlined values indicate second-best results.
Metrics Original BeautyGAN LADN PSGAN++ Clip2Protect DiffAM MAD
Ours-DiffClean
Ours-DiffClean
Minor/Adult Acc (%) 82.8 86.12 84.34 82.7 83.8 85.5 72.53 87.8 88.6
Age group Acc (%) 34.4 35.57 30.9 32.54 33.3 34.5 28.03 36.8 37.0
MAE 6.56 6.09 6.98 6.95 6.54 6.32 9.51 5.76 5.71
@@
BeautyGAN 72.1 ($\downarrow$) 83.2 ($\uparrow$)
LADN 45.2 ($\downarrow$) 58.3 ($\downarrow$)
PSGAN++ 66.2 ($\downarrow$) 75.3 ($\downarrow$)
@@
MAD 22.8 ($\downarrow$) 33.9 ($\downarrow$)
Ours-DiffClean (SSRNet) 75.8 ($\uparrow$) 84.2 ($\uparrow$)
Ours-DiffClean (CLIP) 75.5 ($\uparrow$) 82.4 ($\uparrow$)
@@
Figure 5: (Top): ROC curve with FaceNet. (Middle): ROC curve with MobileFace. (Bottom): Biometric matching in terms of TMR (%) @FMR = 0.01% (higher is better) with FaceNet and MobileFace matchers on FFHQ. Bolded values indicate best results while underlined values indicate second-best results.
V-B Identity verification
@@
We present the results of biometric identity verification using FaceNet and MobileFace face recognition (FR) models in terms of True Match Rate (TMR) at a False Match Rate FMR=0.01%; see Fig. 5. First, we compute the Original performance by comparing the clean FFHQ face images (without makeup) with images simulated using EleGANt (with makeup). Original performance indicates that makeup degrades biometric matching compared to clean images. Next, we present the performances of the baseline algorithms and DiffClean by comparing the clean FFHQ face images (without makeup) with the outputs of the respective makeup removal algorithms (makeup removed images). It is important to note that most baselines either reported human evaluations (e.g., LADN and BeautyGAN), or only reported genuine scores (e.g., PSGAN++ and MAD), while adversarial makeup transfer methods (e.g., CLIP2Protect and DiffAM) reported the attack success rate. In contrast, we perform both genuine comparisons to indicate successful makeup removal and imposter comparisons to ensure that artifacts should not cause spurious false matches. We observe that baseline performances are lower or comparable with Original indicating higher false matches due to spurious artifacts or unsuccessful makeup removal. On the other hand, DiffClean improves the True Match Rate (TMR) by 1.3% with FaceNet and 5.1% with MobileFace compared to Original performance.
Figure 6: Results of DiffClean on real-world makeup images from LADN [ladn_2019] (left), BeautyFace [Beautyface] (center), and Makeup- Wild [jiang2020psgan] (right) datasets.
Figure 7: Examples from Makeup-Wild dataset with variations in pose, illumination, background and facial accessories. DiffClean preserves details such as fingers, jewelry, and expressions.
V-C Image quality evaluation
We compared the images before and after makeup removal in terms of image quality metrics: SSIM and PSNR and reported our findings in Table II. We compute the Original performance by comparing the clean FFHQ face images (without makeup) with images simulated using EleGANt (with makeup). Our algorithm is at par with DiffAM in terms of SSIM=0.98 while being second best in terms of PSNR=35.64, showing high retention of visual fidelity.
[…]
SSIM PSNR
Original 0.58 14.94
BeautyGAN 0.89 21.48
LADN 0.43 13.53
PSGAN++ 0.85 20.65
CLIP2Protect 0.46 15.74
DiffAM 0.98 39.52
MAD 0.92 26.46
Ours-DiffClean (SSRNet)
0.98 35.64
Ours-DiffClean (CLIP)
0.98 35.29
V-D Real-world generalizability
Real-world use cases, such as, online age verification only include a facial image with or without makeup, it does not include the individual’s actual age label. So, in this experiment, we applied DiffClean to real-world makeup images from the BeautyFace [Beautyface], LADN [ladn_2019], and Makeup-Wild [jiang2020psgan] datasets, and used MiVOLO to predict age before and after makeup removal. See examples of real-world makeup removal in Figs. 6 and 7. In the absence of actual or ground truth age labels, we must rely on statistical measures based on overestimation and underestimation errors to evaluate our performance on real-world datasets; see Table III. Further, we analyzed the confidence interval (CI) using a statistically significant sample size of 3K images from BeautyFace using the student’s t-distribution at a 95% confidence level. We obtain CI = [2.88, 3.09] with margin of error = 0.10 on the underestimation errors, and CI = [2.75, 2.89] with margin of error of 0.07 on the overestimation errors. We also report the identity verification on the LADN dataset using FaceNet and MobileFace matchers in Fig. 8. Both qualitative and quantitative results support that DiffClean is capable of handling various real-world makeup styles with significant diversity in pose and illumination variations.
The robustness of DiffClean across different makeup styles with superior performance in terms of age (Table III) and identity verification (Fig. 8) coupled with its fast inference (2 secs.) makes it well-suited for practical deployment. DiffClean is faster than CLIP2Protect (30 secs.), MAD (3 mins.), and comparable with PSGAN++ (2.35 secs.) and DiffAM (2 secs.).
@@ V-D Real-world generalizability
TABLE III: Statistics of predicted ages on the real-world makeup images from the BeautyFace, LADN and Makeup-Wild datasets before and after makeup removal, and associated overestimation and under estimation errors in predicted age.
@@
After makeup removal
@@
24.64 $\pm$ 3.88 [LADN]
@@
25.56 $\pm$ 4.12 [Makeup-Wild]
DiffClean
24.16 $\pm$ 4.38 2.45 $\pm$ 2.31 3.34 $\pm$ 3.06
DiffClean
24.25 $\pm$ 4.40 2.98 $\pm$ 2.80 3.20 $\pm$ 3.07
FaceNet
MobileFace
Method FR model
FaceNet MobileFace
Ours-DiffClean (SSRNet) 81.6 69.0
Ours-DiffClean (CLIP) 80.2 41.5
Figure 8: (Top): ROC curve with FaceNet. (Middle): ROC curve with MobileFace. (Bottom): Biometric matching in terms of TMR (%) @FMR = 0.01% (higher is better) with FaceNet and MobileFace matchers on LADN dataset.
V-E Further Analysis
V-E1 Training Dataset Diversity
We fine-tuned on 600 images (300 from MT dataset with smoky-eyes, flashy, retro, Korean and Japanese makeup styles + 300 from UTKFace with EleGANt-based makeup transfer with five reference styles and adaptive interpolated shading). To further assess training data diversity, we augmented our training set, $D_{tr}$, and report MAE on validation set, as follows: 4.42 ($\lvert D_{tr}\rvert=600$), 4.55 ($\lvert D_{tr}\rvert=1,200$) and 4.68 ($\lvert D_{tr}\rvert=1,600$). We observe consistent MAE across varying training data scale.
V-E2 Influence of each loss component
Our loss is a combination of several losses, each serving a unique role for successful makeup removal: CLIP loss detects presence of makeup. Identity loss restores biometric markers. LPIPS loss avoids perceptual anomaly. Age loss restores facial age characteristics. We report MAE ($\downarrow$) on the validation set by fine-tuning separate models with loss components added incrementally, as follows: 9.28 ($\mathcal{L}_{clip}$), 7.43 ( $\mathcal{L}_{clip}+\mathcal{L}_{id}$), 4.70 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}$), and 4.42 ($\mathcal{L}_{clip}+\mathcal{L}_{id}+\mathcal{L}_{pips}+\mathcal{L}_{SSRNet-age}$). Evidently, the combined losses improve the overall performance.
V-E3 Comparison of $\mathcal{L}_{WSL}$ with $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and Huber loss functions for age estimator
Eqn.(1) refers to the $\mathcal{L}_{WSL}$ which is self-adjusted smoothed $\mathcal{L}_{1}$ loss function adopted in RetinaMask [fu2019retinamask] which is further weighted by a scalar value of 3.0 for the vulnerable age groups between 10-29 yrs. It produced the lowest MAE compared to other losses; see Table IV.
TABLE IV: Comparison of $\mathcal{L}_{WSL}$ with $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and Huber loss functions for age estimator.
MAE with different age losses
Age Group Pretrained L1 Loss MSE Loss Huber Loss $\mathcal{L}_{WSL}$ (Eqn.1)
0-2 20.2 2.1 3.1 2.6 1.6
3-6 27.9 2.3 2.6 2.5 2.2
7-9 29.2 3.7 4.1 3.7 3.1
10-14 26.4 5.0 5.1 5.1 4.6
15-19 18.8 5.6 6.1 6.3 5.3
20-29 14.6 6.4 6.2 6.5 6.1
30-39 12.6 6.6 6.0 6.4 6.0
40-49 9.7 8.0 7.0 7.7 6.8
50-69 9.4 9.7 8.7 8.9 6.9
Average 18.7 5.5 5.4 5.5 4.7
V-E4 Demographic fairness analysis
Table LABEL:Tab:fair reports demographic fairness analysis on FFHQ using DeepFace library [serengil2024deepface], indicating DiffClean does not exhibit demographic disparity. FFHQ has more White images than Black and Indian images, but a balanced gender distribution.
TABLE V: Demographic fairness analysis on FFHQ using our DiffClean SSRNet/CLIP-age loss models in terms of MAE. Values in parentheses indicate count after being processed by DeepFace.